OpenStudy (anonymous):

write in standard form. determine the value of a,b,c x^2+13x-4=14x+1 a= b= c= Give me some solution on how to solve this problem. please

OpenStudy (anonymous):

\[\huge x^2+13x-4=14x+1 \] \[\huge \rightarrow x^2+13x-4-14x-1 =0\] \[\huge \rightarrow x^2-x-5 =0\] is the standard form; comparing it with \[\huge \rightarrow ax^2+bx+c =0\] we find: a= 1, b= -1, c= -5

OpenStudy (anonymous):

@candyhuebner

OpenStudy (anonymous):

thank you for helping me